Mir Status Report 9 April 1997

Date: 9 April 1997 / Mission Day: 88

Experiment Operations: Victor Blagov informed us Energia management has set a deadline of the 15th of April to successfully complete the maintenance activities. If these activities are not finished, then the EVA may be in jeopardy. Furthermore, he requested we clear Jerry's calendar of science through Friday, 11 April, to permit him to work on the maintenance activity. As a result, Jerry will perform only the Sleep experiment and CSST on Friday, April 11. Jerry on a later pass said he does not have any problem performing system maintenance activities, but wanted to make sure the cancellation/delay of science is analyzed carefully.

SMP/Medical: The time reserved for a PMC was used for a regular US comm pass on the open air-to-ground. Soloviev spoke with the crew on the open air-to-ground. He made two 'commands'. First priority: the crew obey the work-rest schedule with a full eight hours of sleep each night. Second priority: No science until station repair operations are complete.

Mir Systems

1. Telemetry - The "MOTsA" unit failed yesterday; this device stores uplinked executable commands, in particular, the command for the MIPS downlink. That means there will be no MIPS downlink available until the system is repaired. Repair activities on this unit are scheduled for 14 April. Suggestion will be made to Jerry to use the Packet system.

2. Repair plan - Two crew members working on removing the leaking heat exchanger from the VGK loop in Kvant 1. The crew will cut and cap the pipes leading in and out of the leaking heat exchanger in the Kvant 1. After the loop has been leak checked, the Vozdukh system will be reactivated. The goal is to complete this activity and reactivate the loop by the end of the 10th.

3. The ground personnel are confident that the remaining three heat exchangers will be adequate to handle the system load. (Late update - The crew was unable to find the IFM kit in the Progress until late this evening. They were told to get organized and begin in the morning.)

4. The other crew member is working on the Elektron system in Kvant 2. The original goal was to repair and re-start the unit on the 9th (hardware for Elektron repair was also only just found late this evening. Elektron work will begin tomorrow as well).

5. Leaks - Leak isolation and detection on the KOB 1 & 2 loops in the Core Module will take four days. The preliminary plan is for this to occur from the 11th through the 14th, although the 12th is a major holiday (Cosmonauts Day), so the schedule is likely to change. A total of 11 leaks have been identified in the KOB -1 and -2 loops in the base block, and the heat exchanger for the VGK in Kvant. None of these loops are operational. In their search for the leak, the crew reported that 'the quality of many of the pipes is not good'. The ground estimates that a total of 3.8 liters of coolant has leaked out this mission.

Environmental Parameters (Information from 7 April)

Parameter                                       Min.      Max.     Range  
Pressure (mmHg) Base Block 749 758 660 - 850
ppO2 (mmHg) 160 167 140 - 200
ppCO2 (mmHg) 5.7 6.4 up to 8.0
ppH20 (mmHg) 15.5 16.1 8 - 14
Humidity (percentage) 53 56 30 - 75
Temperature (0C) base block (BB) 28.2 29.1 18 - 28
Temp (0C) docking/transfer module 22.6 23.1 18 - 28
Temp (0C) work area Krystall 24.7 25.5 18 - 28
Temp (0C) sleep area Krystall 25.8 26.5 18 - 28
Temp (0C) Spektr 26.2 26.9 18 - 28
Temp (0C) Kvant II 26.2 26.9 18 - 28

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on April 9, 1997 9:15 PM.

Mir Status Report 8 April 1997 was the previous entry in this blog.

Shuttle Mir Science Program Status Report 3-9 April 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.